The weight of an object varies inversely with the square of the distance from earth's center (3,960 miles). A man weighs 250 lbs on the surface of the earth. How much would he weigh in an airplane flying at 30,000 ft?: The weight of an object varies inversely with the square of the distance from earth's center (3,960 miles). A man weighs 250 lbs on the surface of the earth. How much would he weigh in an airplane flying at 30,000 ft?
Answer by Alan3354(1926)

The weight of an object varies inversely with the square of the distance from earth's center (3,960 miles). A man weighs 250 lbs on the surface of the earth. How much would he weigh in an airplane flying at 30,000 ft?
------------------------------
3960 miles = 3960*5280 feet = 20908800 feet
His weight at 30000 feet above that is
250*((20908800)/(20908800+30000))^2
= 250*(209088/209388)^2
= 249.284 pounds
That's a difference of 0.716 pounds, or 11.45 ounces.